OpenStudy (anonymous):

lim t->0 [1/(t√(1+t))-(1/t)]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 16} (\frac{ 1 }{ t \sqrt{1+t} } - \frac{ 1 }{ t })\]

OpenStudy (dinamix):

(1-(17)^1/2 /(17)^1/2) 1/16

OpenStudy (michele_laino):

hint: your function, is continue at t=16

OpenStudy (anonymous):

*as t approaches 0... sorry

OpenStudy (michele_laino):

you have to replace t with 16 only

OpenStudy (dinamix):

its \[-\]

OpenStudy (dinamix):

when t---->0

OpenStudy (anonymous):

the answer is -1/2

OpenStudy (anonymous):

yes, the answer is -1/2 @vvvbb

OpenStudy (anonymous):

yes... so how do I get there?

OpenStudy (anonymous):

 choose common denominator and simplify

OpenStudy (anonymous):

then you get\[\frac{ 1-\sqrt{1+t} }{ t \sqrt{1+t}}\]

OpenStudy (anonymous):